David likes playing cricket. This week he scored seven more than four times as many runs as last week. if he scored 67 runs this week, how many runs did David score last week?
Answer by ptaylor(2198)


Let x= number of runs scored last week
4x+7=seven more than four times as many runs as last week
So, our equation to solve is:
4x+7=67 subtract 7 from each side
4x+7-7=67-7 collect like terms
4x=60 divide each side by 4
x=15----------------number of runs scored last week
CK
4x+7=4*15+7=67 ---number of runs scored this week

Hope this helps---ptaylor